# Vector Field-Guided Learning Predictive Control for Motion Planning of Mobile Robots with Unknown Dynamics

### v1

@@ IV-A1 Composite Vector Field with Kinodynamic Constraints
$\chi_{c}(\xi)=$ $\left(\prod_{i\in\mathcal{I}\backslash\mathcal{I}^{\prime}}\sqcup_{\mathcal{Q}_{i}}(\xi)\prod_{i\in\mathcal{I}^{\prime}}s_{i}(\xi)\right)\hat{\chi}_{\mathcal{P}}(\xi)$ (13)
$+\sum_{i\in\mathcal{I}\backslash\mathcal{I}^{\prime}}\left(\sqcap_{\mathcal{R}_{i}}(\xi)\hat{\chi}_{\mathcal{R}_{i}}(\xi)\right)+\sum_{i\in\mathcal{I}^{\prime}}\left(1-{s_{i}}(\xi)\right)\hat{\chi}_{\mathcal{R}_{i}}(\xi),$
@@
The placed virtual obstacle is assumed to satisfy $({}^{\mathrm{in}}\mathcal{Q}_{\rm{vrtl}}\cap\mathcal{P})\subset({}^{\mathrm{in}}\mathcal{R}_{\rm{actl}}\cap\mathcal{P})$. The assumption is used for letting $\xi$ exit the sandwiched region $\mathcal{M}_{s}$ from $\mathcal{M}_{b}$, but not converging to $\mathcal{Q}_{\rm{vrtl}}$ when robots are in $\mathcal{M}_{b}$.
@@
Virtual obstacles cease to exert their influence if $\xi$ enter $\mathcal{M}_{b}$ again from $\mathcal{M}_{s}$. This setting stops robots from returning to $\mathcal{M}_{s}$, but enables them to move towards the desired path to complete an obstacle avoidance process.
@@ IV-B Online Compensation to Update the Offline Trained Deep Koopman Model
Due to the necessity for efficient computation in the online compensation of vehicle dynamic models, we further apply the approximate linear dependence (ALD) strategy[9] to quantize the online updated dictionary $\mathcal{D}_{\rm{GP}}$ for obtaining the training dataset $\mathcal{D}_{\rm{SGP}}$. Note that $\mathcal{D}_{\rm{GP}}$ is obtained in the same way as [36]. Subsequently, we employ the Sparse GP-based algorithm FITC [33] for online estimation to enhance the computational efficiency. This approach aligns with our previous work, and additional details can be found in[9].
@@
where $A_{d,k}\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$, $B_{d,k}\in\mathbb{R}^{n_{\mathcal{K}}\times n_{u}}$. Note that $A_{d,k}$ and $B_{d,k}$ are then used for the training process of model-based RL in Section IV-D. To obtain $\frac{\partial{d}(z_{k})}{\partial\varUpsilon{{x}_{k}})}$ and $\frac{\partial{d}(z_{k})}{\partial{{u}_{k}}}$, one should compute $\frac{\partial{d}(z_{k})}{\partial{{z}_{k}}}$ first, which is equivalent to calculating $\partial\left(Q_{z_{k}\mathbf{z}_{\rm{sp}}}(Q_{\mathbf{z}_{\rm{sp}}\mathbf{z}_{\rm{sp}}}+\Lambda)^{-1}\mathbf{y}_{\rm{sp}}\right)/\partial z_{k}$, i.e.,
@@ IV-C Exponential Barrier Function Incorporated Cost Function to Avoid Moving Obstacles
Due to the adoption of rule-based switching between tracking control and obstacle-avoiding modes in [9], without considering obstacle velocities, it does not guarantee the safety of mobile robots under extreme conditions such as the hazardous driving behavior of opposing vehicles. Our idea stems from the work[37] in completing suicidal game tasks. In [37], the “evader” aspires to avoid being caught by the “pursuer” by computing safe actions. Similarly, when moving obstacles obstruct the road ahead, our strategy is to keep a safe distance from obstacles and track the desired trajectory if the ego robot cannot be caught. Integrating [9] and [37], the exponential barrier function comes into effect if the nearest surrounding obstacle would collide with the robot, where the collision judgment is generated by the method in [37], as elaborated below.
@@ IV-D Online Receding-horizon Reinforcement Learning with a Game-based Exponential Barrier Function Design
To minimize the infinite-horizon cost function $V_{\infty}(\tilde{x}_{k})$, we first present the definitions of the optimal control and value functions commonly used in infinite-horizon ADP and RL. As a class of ADP methods, the dual heuristic programming (DHP) approach aims to minimize the derivative of the value function with respect to the state. Consistent with the DHP-based framework, we re-define the value function as follows:
@@
Then, the optimal control corresponding to the infinite-horizon optimization objective is defined by the following equation:
$\tilde{u}_{k}^{*}=\arg\min_{\tilde{u}_{k}}\sum_{\tau=k}^{\infty}\gamma^{\tau-k}L(\tilde{x}_{\tau},\tilde{u}_{\tau}),$ (27)
@@
Note that the optimal control and value function can be solved under the iterative learning framework[38, 9]. However, the online convergence property is subject to the learning efficiency. Motivated by the concept of receding-horizon from MPC, as discussed in [39], we divide $V_{\infty}(\tilde{x}_{k})$ into multiple sub-problems within the prediction horizon $[k,k+N]$ to compute an admissible control $\tilde{u}_{\tau}$ by optimizing $V(\tilde{x}_{\tau})$, where $N$ is the prediction horizon length. For any time instant $\tau$ in the prediction horizon, we have
@@
where $\mathbb{E}[\cdot]$ denotes the expectation value, and $F(\hat{\tilde{x}}_{k+N})$ is the terminal cost function, defined by $F(\hat{\tilde{x}}_{k+N})=\hat{\tilde{x}}_{k+N}^{\top}P\hat{\tilde{x}}_{k+N}$, where $P\succ 0\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$ is the terminal penalty matrix, which can be determined in the same way as [40].
@@
$V^{i+1}(\tilde{x}_{\tau})=\mathbb{E}\left[L\left(\tilde{x}_{\tau},\tilde{u}_{\tau}\right)+\gamma V^{i}(\hat{\tilde{x}}_{\tau+1})\right],\tau\in[k,k+N-1],$ (29)
@@
$\lambda(\tilde{x}_{\tau})=\partial V(\tilde{x}_{\tau})/\partial\tilde{x}_{\tau}.$ (30)
@@
For the critic network, the optimal value function corresponding to the finite-horizon optimization objective is defined by $\partial V^{*}(\tilde{x}_{\tau})/\partial\tilde{x}_{\tau}$, i.e,
$\lambda^{*}(\tilde{x}_{\tau})=\left\{\begin{array}[]{cc}2Q\tilde{x}_{\tau}+\gamma A_{d,\tau}^{\top}\lambda^{*}(\hat{\tilde{x}}_{\tau+1})+\dfrac{\partial h({x}_{\tau})}{\partial\tilde{x}_{\tau}},&\tau\in[k,k+N-1]\\ 2P\hat{\tilde{x}}_{k+N}+{\partial h({{x}}_{k+N})}/{\partial\hat{\tilde{x}}_{k+N}},&\tau=k+N\end{array}\right.$
With the optimal value function $\lambda^{*}(\tilde{x}_{\tau})$, one can compute the optimal control by setting ${\partial V^{*}(\tilde{x}_{\tau})}/{\partial\tilde{u}(\tilde{x}_{\tau})}=0$. Then, for the actor network, the optimal control corresponding to the finite-horizon optimization objective is defined by the following equation:
$\tilde{u}^{*}(\tilde{x}_{\tau})=u_{b}\otimes\tanh(-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\lambda^{*}(\hat{\tilde{x}}_{\tau+1})),$ (32)
where $u_{b}=[u_{1},\cdots,u_{n_{u}}]^{\top}\in\mathbb{R}^{n_{u}}$ and $u_{1},\cdots,u_{n_{u}}$ denote the maximum values of the corresponding control dimensions, and $\tanh(\cdot):\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}^{n_{u}}$ is the hyperbolic tangent function.
Input: Obstacle states, the ego robot’s states, the preliminary reference path.
@@
Output: The optimal control sequence $\{\tilde{u}^{*}_{1},\tilde{u}^{*}_{2},\cdots\}$.
@@ 4 Function LPC:
15                  Compute $\hat{\tilde{u}}(\tilde{x}_{\tau})$ and $\hat{\lambda}(\tilde{x}_{\tau})$ by Eq. (33).
@@
17                  Compute $\epsilon_{c}^{i}(\tilde{x}_{\tau})$ and $\epsilon_{a}^{i}(\tilde{x}_{\tau})$ with (36) and (37).
@@
19                  Apply $\hat{\tilde{u}}^{i}(\tilde{x}_{\tau})$ to the system (11).
@@
23                   Obtain the current control policy, and exit.
@@
      // Control Horizon
27       Compute the current control $\hat{\tilde{u}}^{i}(\tilde{x}_{\tau})$ with (33a).
@@ 37 Function Main():
42
43                  Obtain the desired trajectory with Eqs. (13) and (17).
@@
49
50            Perform LPC and apply $\hat{\tilde{u}}(\tilde{x}_{\tau})$ to the robot.
@@
The optimal control $\tilde{u}^{*}(\tilde{x}_{\tau})$ and the optimal value function $\lambda^{*}(\tilde{x}_{\tau})$ are difficult to be solved analytically. This is because $\lambda^{*}(\hat{\tilde{x}}_{\tau+1})$ is not available; see Eqs. (31) and (32). However, they can be approximated by two kernel-based structures, which are
$\hat{\tilde{u}}(\tilde{x}_{\tau})={W}_{a,\tau}^{\top}\Phi(\tilde{x}_{\tau}),$ (33a)
$\hat{\lambda}(\tilde{x}_{\tau})={W}_{c,\tau}^{\top}\Phi(\tilde{x}_{\tau}),$ (33b)
where ${W}_{a,\tau}\in\mathbb{R}^{n_{\Phi}\times n_{u}}$ and ${W}_{c,\tau}\in\mathbb{R}^{n_{\Phi}\times n_{\mathcal{K}}}$ are the weights of the actor and critic networks, respectively, and $n_{\Phi}$ denotes the dimension of the RL training dictionary, which is obtained by employing the ALD strategy[9] to select representative samples from the robot’s state space. For the current state $\tilde{x}_{k}$, the basis function in Eq. (33) is constructed by
$\Phi(\tilde{x}_{\tau})=\left[k(\tilde{x}_{\tau},c_{1}),\ldots,k(\tilde{x}_{\tau},c_{n_{\Phi}})\right]^{\top}\in\mathbb{R}^{n_{\Phi}},$
@@
With the estimated value function in (33b) and consistent with (32), we can obtain the target optimal control at the $\tau$-th horizon by
$\tilde{u}(\tilde{x}_{\tau})=u_{b}\tanh\left(-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\hat{\lambda}(\hat{\tilde{x}}_{\tau+1})\right).$ (34)
@@
$\lambda(\tilde{x}_{\tau})=\left\{\begin{array}[]{cc}2Q\tilde{x}_{\tau}+\gamma A_{d,\tau}^{\top}\hat{\lambda}(\hat{\tilde{x}}_{\tau+1})+\dfrac{\partial h({x}_{\tau})}{\partial\tilde{x}_{\tau}},&\tau\in[k,k+N-1]\\ 2P\hat{\tilde{x}}_{k+N}+{\partial h({{x}}_{k+N})}/{\partial\hat{\tilde{x}}_{k+N}},&\tau=k+N\end{array}\right.$
Next, we present the iterative learning processes. From the previous formulations, $N$ actor-critic networks are to be constructed to approximate the optimal value function and control in (31) and (32). At the $i$-th iteration, the critic network aims at minimizing the error function between the target and the approximate value functions, which is defined by
$\epsilon_{c}^{i}(\tilde{x}_{\tau})=\frac{1}{2}\lVert\lambda^{i}(\tilde{x}_{\tau})-\hat{\lambda}^{i}(\tilde{x}_{\tau})\rVert^{2}.$ (36)
For the actor network, it minimizes the error function between the target and the approximate controls, which is defined by the following equation:
$\epsilon_{a}^{i}(\tilde{x}_{\tau})=\frac{1}{2}\lVert\tilde{u}^{i}(\tilde{x}_{\tau})-\hat{\tilde{u}}^{i}(\tilde{x}_{\tau})\rVert^{2}.$ (37)
@@
${W}_{c,\tau+1}^{i}$ $={W}_{c,\tau}^{i}-\eta_{c}{\partial\epsilon_{c}^{i}(\tilde{x}_{\tau})}/{\partial W_{c,\tau}^{i}},$ (38a)
${W}_{a,\tau+1}^{i}$ $={W}_{a,\tau}^{i}-\eta_{a}{\partial\epsilon_{a}^{i}(\tilde{x}_{\tau})}/{\partial W_{a,\tau}^{i}},$ (38b)
@@
Theorem 1 indicates that the Eb-RHRL approach in Algorithm 1 solves the optimal safe control problems under state and control constraints by obtaining the near-optimal solutions.

### v2

@@ IV-A1 Composite Vector Field with Kinodynamic Constraints
[…]
@@
[…]
@@
[…]
@@ IV-B Online Compensation to Update the Offline Trained Deep Koopman Model
[…]
@@
where $A_{d,k}\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$, $B_{d,k}\in\mathbb{R}^{n_{\mathcal{K}}\times n_{u}}$. Note that $A_{d,k}$ and $B_{d,k}$ are then used for the training process of model-based RL in Section IV-D. To obtain $\frac{\partial{d}(z_{k})}{\partial\varUpsilon({{x}_{k}})}$ and $\frac{\partial{d}(z_{k})}{\partial{{u}_{k}}}$, one should compute $\frac{\partial{d}(z_{k})}{\partial{{z}_{k}}}$ first, which is equivalent to calculating $\partial\left(Q_{z_{k}\mathbf{z}_{\rm{sp}}}(Q_{\mathbf{z}_{\rm{sp}}\mathbf{z}_{\rm{sp}}}+\Lambda)^{-1}\mathbf{y}_{\rm{sp}}\right)/\partial z_{k}$, i.e.,
@@ IV-C Exponential Barrier Function Incorporated Cost Function to Avoid Moving Obstacles
Due to the adoption of rule-based switching between tracking control and obstacle-avoiding modes in [9], without considering obstacle velocities, it does not guarantee the safety of mobile robots under extreme conditions such as the hazardous driving behavior of opposing vehicles. Our idea stems from the work[37] in completing suicidal game tasks, where the “evader” aspires to avoid being caught by the “pursuer” by computing safe actions. Similarly, when moving obstacles obstruct the road ahead, our strategy is to keep a safe distance from obstacles and track the desired trajectory if the ego robot cannot be caught. Integrating [9] and [37], the exponential barrier function comes into effect if the nearest surrounding obstacle would collide with the robot, where the collision judgment is generated by the method in [37], as elaborated below.
@@ IV-D Online Receding-horizon Reinforcement Learning with a Game-based Exponential Barrier Function Design
To minimize the infinite-horizon value function $V_{\infty}(\tilde{x}_{k})$, we first present the definitions of the optimal solution (control) and value functions commonly used in infinite-horizon ADP and RL. As a class of ADP methods, the dual heuristic programming (DHP) approach aims to minimize the derivative of the value function with respect to the state. Consistent with the DHP-based framework, we re-define the value function as
@@
Then, the optimal solution corresponding to the infinite-horizon optimization objective is defined by the following equation:
$\tilde{u}_{k}^{*}=\arg\min_{\tilde{u}_{k}}\sum_{\tau=k}^{\infty}\gamma^{\tau-k}L(\hat{\tilde{x}}_{\tau},\tilde{u}_{\tau}),$ (27)
@@
Note that the optimal solution and value function can be solved under the iterative learning framework[9]. However, the online convergence property is subject to the learning efficiency. Motivated by the concept of receding-horizon from MPC, as discussed in [11], we divide $V_{\infty}(\tilde{x}_{k})$ into multiple sub-problems within the prediction horizon $[k,k+N]$, where $N$ is the prediction horizon length. For any time instant $\tau$ in the prediction horizon, we have
@@
where $\mathbb{E}[\cdot]$ denotes the expectation value, and $F(\hat{\tilde{x}}_{k+N})$ is the terminal cost function, defined by $F(\hat{\tilde{x}}_{k+N})=\hat{\tilde{x}}_{k+N}^{\top}P\hat{\tilde{x}}_{k+N}$, where $P\in\mathbb{R}^{n_{\mathcal{K}}\times n_{\mathcal{K}}}$ is the terminal penalty matrix, which can be determined in the same way as [11].
@@
$V^{i+1}(\hat{\tilde{x}}_{\tau})=\mathbb{E}\left[L(\hat{\tilde{x}}_{\tau},\tilde{u}_{\tau})+\gamma V^{i}(\hat{\tilde{x}}_{\tau+1})\right],\tau\in[k,k+N-1],$ (29)
@@
$\lambda(\hat{\tilde{x}}_{\tau})=\partial V(\hat{\tilde{x}}_{\tau})/\partial\hat{\tilde{x}}_{\tau}.$ (30)
@@
For the critic network, the optimal value function corresponding to the finite-horizon optimization objective is defined by $\partial V^{*}(\hat{\tilde{x}}_{\tau})/\partial\hat{\tilde{x}}_{\tau}$, i.e,
$\lambda^{*}(\hat{\tilde{x}}_{\tau})=\left\{\begin{array}[]{cc}2Q\hat{\tilde{x}}_{\tau}+\gamma A_{d,\tau}^{\top}\lambda^{*}(\hat{\tilde{x}}_{\tau+1})+\dfrac{\partial h(\hat{x}_{\tau})}{\partial\hat{\tilde{x}}_{\tau}},&\tau\in[k,k+N-1]\\ 2P\hat{\tilde{x}}_{k+N}+{\partial h({\hat{x}}_{k+N})}/{\partial\hat{\tilde{x}}_{k+N}},&\tau=k+N\end{array}\right.$
With the optimal value function $\lambda^{*}(\hat{\tilde{x}}_{\tau})$, one can compute the optimal solution by setting ${\partial V^{*}(\hat{\tilde{x}}_{\tau})}/{\partial\tilde{u}(\hat{\tilde{x}}_{\tau})}=0$. Then, for the actor network, the optimal solution corresponding to the finite-horizon optimization objective is defined by the following equation:
$\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})=-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\lambda^{*}(\hat{\tilde{x}}_{\tau+1}).$ (32)
Input: Obstacle information, robot’s states, the desired path.
@@
Output: The final safe trajecotry.
@@ 4 Function LPC:
15                  Compute $\hat{\tilde{u}}^{i}(\hat{\tilde{x}}_{\tau})$ and $\hat{\lambda}^{i}(\hat{\tilde{x}}_{\tau})$ by Eq. (33).
@@
17                  Compute $\epsilon_{c}^{i}(\hat{\tilde{x}}_{\tau})$ and $\epsilon_{a}^{i}(\hat{\tilde{x}}_{\tau})$ with (36) and (37).
@@
19                  Apply $\hat{\tilde{u}}^{i}(\hat{\tilde{x}}_{\tau})$ to the system (11).
@@
23                   ${W}_{c,k:k+N}^{0}\leftarrow W_{c,k:k+N}^{i}$, ${W}_{a,k:k+N}^{0}\leftarrow W_{a,k:k+N}^{i}$.
24                  Return the sequence $\{\hat{\tilde{u}}^{i}(\tilde{x}_{k}),\cdots,\hat{\tilde{u}}^{i}(\hat{\tilde{x}}_{k+N})\}$.
@@ 38 Function Main():
43
44                  Generate the guidance (the desired trajectory) for the LPC module with Eqs. (13) and (17).
@@
50
            // Control Horizon
51             Perform LPC and apply the 1st solution of the sequence to the real-world robot.
@@
The optimal solution $\tilde{u}^{*}(\hat{\tilde{x}}_{\tau})$ and the optimal value function $\lambda^{*}(\hat{\tilde{x}}_{\tau})$ are difficult to be solved analytically. This is because $\lambda^{*}(\hat{\tilde{x}}_{\tau+1})$ is not available; see Eqs. (31) and (32). However, they can be approximated by two kernel-based networks, i.e.,
$\hat{\tilde{u}}(\hat{\tilde{x}}_{\tau})={W}_{a,\tau}^{\top}\Phi(\hat{\tilde{x}}_{\tau}),$ (33a)
$\hat{\lambda}(\hat{\tilde{x}}_{\tau})={W}_{c,\tau}^{\top}\Phi(\hat{\tilde{x}}_{\tau}),$ (33b)
where ${W}_{a,\tau}\in\mathbb{R}^{n_{\Phi}\times n_{u}}$ and ${W}_{c,\tau}\in\mathbb{R}^{n_{\Phi}\times n_{\mathcal{K}}}$ are the weights of the actor and critic networks, respectively, and $n_{\Phi}$ denotes the dimension of the RL training dictionary, which is obtained by employing the ALD strategy[9] to select representative samples from the robot’s state space. For the current state $\hat{\tilde{x}}_{\tau}$, the basis function in Eq. (33) is constructed by
$\Phi(\hat{\tilde{x}}_{\tau})=\left[k(\hat{\tilde{x}}_{\tau},c_{1}),\ldots,k(\hat{\tilde{x}}_{\tau},c_{n_{\Phi}})\right]^{\top}\in\mathbb{R}^{n_{\Phi}},$
@@
With the estimated value function in (33b) and consistent with (32), we can obtain the target optimal solution at the $\tau$-th horizon by
$\tilde{u}(\hat{\tilde{x}}_{\tau})=-\dfrac{1}{2}\gamma R^{-1}B_{d,\tau}^{\top}\hat{\lambda}(\hat{\tilde{x}}_{\tau+1}).$ (34)
@@
$\lambda(\hat{\tilde{x}}_{\tau})=\left\{\begin{array}[]{cc}2Q\hat{\tilde{x}}_{\tau}+\gamma A_{d,\tau}^{\top}\hat{\lambda}(\hat{\tilde{x}}_{\tau+1})+\dfrac{\partial h(\hat{x}_{\tau})}{\partial\hat{\tilde{x}}_{\tau}},&\tau\in[k,k+N-1]\\ 2P\hat{\tilde{x}}_{k+N}+{\partial h({\hat{x}}_{k+N})}/{\partial\hat{\tilde{x}}_{k+N}},&\tau=k+N\end{array}\right.$
Next, we present the iterative learning processes. From the previous formulations, $N$ actor-critic networks are to be constructed to approximate the optimal value function and solution in (31) and (32). At the $i$-th iteration, the critic network aims at minimizing the error function between the target and the approximate value functions, i.e.,
$\epsilon_{c}^{i}(\hat{\tilde{x}}_{\tau})=\frac{1}{2}\lVert\lambda^{i}(\hat{\tilde{x}}_{\tau})-\hat{\lambda}^{i}(\hat{\tilde{x}}_{\tau})\rVert^{2}.$ (36)
For the actor network, it minimizes the error function between the target and the approximate solutions, which is defined by the following equation:
$\epsilon_{a}^{i}(\hat{\tilde{x}}_{\tau})=\frac{1}{2}\lVert\tilde{u}^{i}(\hat{\tilde{x}}_{\tau})-\hat{\tilde{u}}^{i}(\hat{\tilde{x}}_{\tau})\rVert^{2}.$ (37)
@@
${W}_{c,\tau+1}^{i}$ $={W}_{c,\tau}^{i}-\eta_{c}{\partial\epsilon_{c}^{i}(\hat{\tilde{x}}_{\tau})}/{\partial W_{c,\tau}^{i}},$ (38a)
${W}_{a,\tau+1}^{i}$ $={W}_{a,\tau}^{i}-\eta_{a}{\partial\epsilon_{a}^{i}(\hat{\tilde{x}}_{\tau})}/{\partial W_{a,\tau}^{i}},$ (38b)
@@
Theorem 1 indicates that the Eb-RHRL approach in Algorithm 1 solves the optimal IMPC problems under state constraints by obtaining the near-optimal solutions.